OpenStudy (anonymous):

Please Help A sociology seminar has 9 members. They begin each meeting with each member shaking hands with all the other members. No one shakes another member's hand twice. How many handshakes take place

OpenStudy (anonymous):

Let's label all nine members, A,B,C,D,E,F,G,H,I Now "each member shaking hands with all other memeber" simply means everybody shakes hands. So A must shake everybody's hands: AB AC AD AE AF AG AH AI That's 8 handshakes in total. B must shake everybody's hand, and has already shaken A's: BC BD BE BF BG BG BI That's 7+8 handshakes in total. And so on. The formula is \[\sum_{i=1}^{n-1}i\] Here n=9, So TOTAL=1+2+3+4+5+6+7+9=37

OpenStudy (anonymous):

I understand the different combinations that were made. Would I add the number of handshakes with the total of n=9?

OpenStudy (anonymous):

No, why would you? Think about it. The total number of handshakes is all you need.

OpenStudy (anonymous):

Yeah I barely realized that.. sorry for asking that question. Thank You Very Much For your help I really appreciate it =)